What is the vapor pressure of a 1.00 molal sugar solution at 25°C? Sugar is a nonvolatile, nonelectrolyte solute. The vapor pressure of water at 25°C is 23.8 torr.
1
Expert's answer
2020-12-14T14:43:38-0500

[Psaturated(water)-P(water)]/Psaturated(water)=1-xwater=xsugar, x is the mole fraction

1 molal solution => 1 mol of sugar and 1 kg of water => 1 mol of sugar and 1000/18.015=55.402 mol of water => xsugar = (1 mol)/(1 mol + 55.402 mol) = 0.0177

P(solution) = P(water) = - xsugar*Psaturated(water) + Psaturated(water) = - 0.0177*23.8 + 23.8 = 23.38 torr.